Convert each of the following direct quotes to dollar indirect quotes:

1 Danish krone = $0.1775 (Round your answer to 3 decimal places.)

exchange rate:

1 Indian rupee = $0.0225 (Round your answer to 3 decimal places.)

exchange rate:

1 Israeli shekel = $0.2790 (Round your answer to 3 decimal places.)

exchange rate
Answer #1
Case Indirect quoate
1 $ =(1/0.1775)=5.634 Danis krone
1 $ =(1/0.0225)=44.444 indian rupee
1 $ =(1/0.2790)=3.584 Israeli shekel(Approx)